# Federated Cubic Regularized Newton Learning with Sparsification-amplified Differential Privacy

### v1

@@ 4 Privacy Analysis
$\left|\left[\nabla l(x,\zeta_{i}^{(j)})\right]_{h}\right|\leq\frac{G_{1}}{\sqrt{d}},\quad\left\|\left[\nabla^{2}l(x,\zeta_{i}^{(j)})\right]_{h}\right\|\leq\frac{G_{2}}{\sqrt{d}}$
@@
Assumption 4 characterizes the sensitivity of each coordinate of gradient $\nabla l(x,\zeta_{i}^{(j)})$ and each row of Hessian $\nabla^{2}l(x,\zeta_{i}^{(j)})$, and implies $\left\|\nabla l(x,\zeta_{i}^{(j)})\right\|\leq G_{1}$ and $\left\|\nabla^{2}l(x,\zeta_{i}^{(j)})\right\|_{2}\leq\left\|\nabla^{2}l(x,\zeta_{i}^{(j)})\right\|_{F}\leq G_{2}$, which can be enforced by the gradient and Hessian clip techniques [23].
To analyze the interplay between the sparsification and privacy, let $c_{i}^{t}$ denote the randomly selected coordinate set for client $i$ at round $t$, i.e., $\mathcal{S}(\cdot)=\frac{d}{k}[\cdot]_{c_{i}^{t}}$. An important observation is that only the values in $c_{i}^{t}$ are transmitted to the central server, i.e.,
$\mathcal{S}(y_{i,t})=\frac{d}{k}\left[\frac{x_{i,t+1}-x_{i,t}}{\alpha}\right]_{c_{i}^{t}}=\frac{d}{\alpha k}\left([x_{i,t+1}]_{c_{i}^{t}}-[x_{i,t}]_{c_{i}^{t}}\right).$
@@
$\leq$ $\frac{2\sqrt{k}G_{1}}{\sqrt{d}}+\frac{2\sqrt{k}G_{2}D}{\sqrt{d}}.$
Lemma 1 indicates that the required noise intensity for achieving $(\varepsilon,\delta)$-DP relies on the sensitivity value. From (4), sparsification reduces the conventional sensivity $2(G_{1}+G_{2}D)$ by a factor of $\sqrt{k/d}$. Consequently, sparsification decreases sensitivity, leading to a reduction in noise intensity. Theorem 1 states a sufficient condition for achieving $(\varepsilon,\delta)$-DP based on the reduced sensitivity resulting from sparsification.
@@
Suppose that Assumptions 1 and 3 are satisfied, and the random-$k$ sparsifier with $k\leq d$ is used in Algorithm 1. Given parameters $m$, $\tau$, $\varepsilon\in(0,1]$, and $\delta_{0}\in(0,1]$, if
$\sigma^{2}\geq\frac{80\tau Tk\log(1.25/\delta_{0}){(G_{1}+G_{2}D)}^{2}}{\varepsilon^{2}m^{2}d}$ (7)
and $T\geq\frac{\varepsilon^{2}}{4\tau}$, then DP-FCRN is $(\varepsilon,\delta)$-DP given certain constant $\delta\in(0,1]$.
@@
The required noise intensity is proportional to the sparsification ratio, $k/d$. Therefore, to achieve the same level of DP, the required noise under our algorithm can be reduced by decreasing $k$. In other words, the fewer transmitted bits, the less noise required for $(\varepsilon,\delta)$-DP.
@@ 5 Convergence Analysis
This section presents the convergence analysis of Algorithm 1. We begin by introducing some properties of random-$k$ sparsification.
Lemma 2.
In each step of the algorithm, a global cubic upper bound function $\phi:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ for $f(w)$ is constructed as
@@
Algorithm 2 uses the typical SGD to solve (5) and the local cubic upper bound $\phi_{i}(x;x_{t})$ is a strongly convex function. Therefore, we can obtain the suboptimality gap based on SGD analysis.
@@
Suppose that Assumptions 13 hold. For every $\beta\in(0,1)$, given parameters $\varepsilon\in(0,1]$, $\delta_{0}\in(0,1]$, and $w\in\mathcal{X}$ the output of Algorithm 2, if we set the number of local iterations as
$\tau=\frac{(L_{0}+L_{1}D+MD^{2}/2)^{2}\varepsilon^{2}m^{2}}{kT\log(1/\delta_{0})(G_{1}+G_{2}D)^{2}},$ (9)
@@
$=$ $O\left(\frac{k\log(1/\delta_{0}){(G_{1}+G_{2}D)}^{2}T}{\varepsilon^{2}m^{2}\mu}\right).$
@@
Suppose that Assumptions 13 hold and the random-$k$ sparsifier with $k\leq d$ is used in Algorithm 1. Then for every $\beta\in(0,1)$, given parameters $m$ and $\varepsilon\in(0,1]$, $\delta_{0}\in(0,1]$, by setting the number of local iterations as (9), the step size as $\alpha>1$ and
$\alpha=O\left(\frac{k\log(1/\delta_{0}){(G_{1}+G_{2}D)}^{2}T}{\varepsilon^{2}m^{2}\mu(L_{0}+L_{1}D+MD^{2}/2)D}\right),$
@@
$\leq$ $\tilde{O}\left(\frac{k\log(1/\delta_{0}){(G_{1}+G_{2}D)}^{2}}{\varepsilon^{2}m^{2}\mu}\cdot\frac{\sqrt{L_{2}}(f(x_{0})-f(x^{*}))^{\frac{1}{4}}}{\mu^{\frac{3}{4}}}\right).$
@@
In existing DP algorithms for strongly convex functions, the best achievable bound for optimization error is $O\left(\frac{d}{\varepsilon^{2}}\right)$ [24, 25]. This indicates that the error bound derived in Theorem 2 is optimal w.r.t. the privacy loss $\varepsilon$. Furthermore, our result $O\left(\frac{k}{\varepsilon^{2}}\right)$ reduces the error bound by a factor $k/d$, attributed to sparsification. This result underscores how efficient communication better balances the trade-off between privacy and utility.
@@
While DP-FCRN does not explicitly include a switching step, the proof of Theorem 2 indicates that DP-FCRN operates in two distinct phases. Initially, when $x_{t}$ is distant from $x^{*}$, the convergence rate is $1/T^{4}$. Subsequently, as $x_{t}$ approaches $x^{*}$, the algorithm transitions to the second phase with a convergence rate of $\exp(\exp(-T))$. In summary, leveraging second-order techniques in our algorithm significantly improves the oracle complexity compared to first-order methods [41].
@@ 6 Numerical Evaluation
In this section, we evaluate the effectiveness of DP-FCRN with different compression levels and compare them to the first-order Fed-SGD with DP [42].
@@ 6.1 Experimental Setup
We test our algorithm on the benchmark datasets epsilon [43], which include 400,000 samples and 2,000 features for each sample. The data samples are evenly and randomly allocated among the $n=40$ clients. The clients cooperatively solve the following logistic regression problem:
@@
As DP parameters, we consider $\varepsilon\in\{0.4,0.6,0.8,1\}$ and $\delta_{0}=0.01$. The random noise is generated according to (7) and the number of local iterations $\tau$ is determined by (9). In iteration $t$, client $i$ processes one data point from $\zeta_{i}$ and the server updates $x_{t}$ accordingly. Upon finishing processing the entire dataset, one epoch is completed. We conduct the algorithm for four epochs and repeat each experiment five times. We show the mean curve along with the region representing one standard deviation. The convergence performance of the algorithm is evaluated by training suboptimality and testing accuracy over iterations. Training suboptimality is calculated by $f(x_{t})-f(x^{*})$, where $f(x^{*})$ is obtained using the LogisticSGD optimizer from scikit-learn [44]. Testing accuracy is determined by applying the logistic function to the entire dataset. It is calculated as the percentage of correct predictions out of the total number of predictions.
@@ 6.2 Performance and Comparison with Fed-SGD
Moreover, employing a larger sparsification ratio $k/d$ in DP-FCRN results in worse training suboptimality, verifying Theorem 2. We find that keeping more coordinates in sparsification leads to more complete information transmission together with increased noise. The results shown in Fig. 1 indicate that, in certain settings, the benefit of noise reduction for convergence performance may outweigh the negative impacts arising from information completeness. On the other hand, there is no obvious difference in testing accuracy with different sparse ratios, which indicates that the performance under the proposed DP-FCRN does not deteriorate much while reducing the communication burden.
@@ 7 Conclusion and Future Work
There are several promising directions for future research. Firstly, investigating methods to reduce the computational complexity of federated second-order learning approaches is valuable. Additionally, integrating more general compression schemes and studying the privacy preservation and performance of non-convex cost functions are intriguing topics.
@@ Appendix A Supporting Lemmas
The following lemma provide some useful properties of $\phi_{i}(v;w)$ [23].
@@
[23] Let $b_{0}>0$ and define the sequence $a_{t+1}\leq b_{0}+\frac{1}{2}a_{t}^{\frac{3}{2}}$ where $a_{0}\leq\frac{16}{9}$. Then, after $T=\Theta(\log\log(\frac{1}{b}))$, we have $a_{T}=O(b_{0})$.
@@ Appendix B Proof of Theorem 1
(Privacy for Subsampling [45]) Suppose $\mathcal{G}$ is an $(\varepsilon,\delta)$-DP mechanism. Consider $\text{Sample}_{r_{1},r_{2}}:\mathcal{D}^{r_{1}}\to\mathcal{D}^{r_{2}}$ as the subsampling manipulation. Given a dataset belonging to $\mathcal{D}^{r_{1}}$ as an input, this subsampling manipulation selects a subset of $r_{2}\leq r_{1}$ elements from the input dataset uniformly at random. For the following mechanism
@@
(Composition of DP [45]) If each of $T$ randomized algorithms $\mathcal{A}_{1},\dots,\mathcal{A}_{T}$ is $(\varepsilon_{i},\delta_{i})$-DP with $\varepsilon_{i}\in(0,0.9]$ and $\delta_{i}\in(0,1]$, then $\mathcal{A}$ with $\mathcal{A}(\cdot)=(\mathcal{A}_{1}(\cdot),\dots,\mathcal{A}_{T}(\cdot))$ is $(\tilde{\varepsilon},\tilde{\delta})$-DP with
@@
$\varepsilon_{s}=\frac{2\sqrt{k\log(1.25/\delta_{0})}{(G_{1}+G_{2}D)}}{\sigma\sqrt{d}}$
@@
$\varepsilon_{s}^{\prime 2}\leq\frac{16k\log(1.25/\delta_{0})(L_{0}+L_{1}D)^{2}}{\sigma^{2}m^{2}d}\leq\frac{\varepsilon^{2}}{5\tau T}\leq 0.8.$
Therefore, we have $\varepsilon_{s}\leq 0.9$ and
$\sum_{s=1}^{\tau T}\varepsilon_{s}^{2}\leq\frac{1}{5}\sum_{s=1}^{\tau T}\varepsilon^{2}\leq 1$ (13)
@@
$\tilde{\varepsilon}=\sqrt{\sum_{s=1}^{\tau T}2\varepsilon_{s}^{2}\log\left(e+\frac{\sqrt{\sum_{s=1}^{\tau T}\varepsilon_{s}^{2}}}{\tilde{\delta}}\right)}+\sum_{s=1}^{\tau T}\varepsilon_{s}^{2}$
@@
$\tilde{\varepsilon}=$ $\sqrt{\sum_{s=1}^{\tau T}2\varepsilon_{s}^{2}\log\left(e+\frac{\sqrt{\sum_{s=1}^{\tau T}\varepsilon_{s}^{2}}}{\tilde{\delta}}\right)}+\frac{1}{5}\varepsilon^{2}$
$\leq$ $\sqrt{3\sum_{s=1}^{\tau T}\varepsilon_{s}^{2}}+\frac{1}{5}\varepsilon$

### v2

@@ 4 Privacy Analysis
$\left|\left[\nabla l(x,\zeta_{i}^{(j)})\right]_{h}\right|\leq\frac{{\color[rgb]{0,0,1}L_{0}}}{\sqrt{d}},\quad\left\|\left[\nabla^{2}l(x,\zeta_{i}^{(j)})\right]_{h}\right\|\leq\frac{{\color[rgb]{0,0,1}L_{1}}}{\sqrt{d}}$
@@
Assumption 2 characterizes the sensitivity of each coordinate of the gradient $\nabla l(x,\zeta_{i}^{(j)})$ and each row of the Hessian $\nabla^{2}l(x,\zeta_{i}^{(j)})$. This assumption is crucial for analyzing the interaction between element selection via sparsification and privacy amplification, as discussed in Hu et al. (2023); Chen et al. (2024). It implies that $\left\|\nabla l(x,\zeta_{i}^{(j)})\right\|\leq{\color[rgb]{0,0,1}L_{0}}$ and $\left\|\nabla^{2}l(x,\zeta_{i}^{(j)})\right\|_{2}\leq\left\|\nabla^{2}l(x,\zeta_{i}^{(j)})\right\|_{F}\leq{\color[rgb]{0,0,1}L_{1}}$, which holds under the assumption of a bounded parameter set.
To analyze the interplay between the sparsification and privacy, let $c_{i}^{t}$ denote the randomly selected coordinate set for client $i$ at round $t$, i.e., $\mathcal{S}(\cdot)=\frac{d}{k}[\cdot]_{c_{i}^{t}}$.
An important observation is that only the values in $c_{i}^{t}$ are transmitted to the central server, i.e.,
${\color[rgb]{0,0,1}\mathcal{S}(y_{i,t})=\frac{d}{k}\left[\alpha(x_{i,t+1}-x_{i,t})\right]_{c_{i}^{t}}=\frac{\alpha d}{k}\left([x_{i,t+1}]_{c_{i}^{t}}-[x_{i,t}]_{c_{i}^{t}}\right).}$
@@
$\leq$ $\frac{2\sqrt{k}{\color[rgb]{0,0,1}L_{0}}}{\sqrt{d}}+\frac{2\sqrt{k}{\color[rgb]{0,0,1}L_{1}}D}{\sqrt{d}}.$
Lemma 1 indicates that the noise intensity required to achieve $(\varepsilon,\delta)$-DP relies on the sensitivity. From (4), sparsification reduces the conventional sensitivity $2({\color[rgb]{0,0,1}L_{0}}+{\color[rgb]{0,0,1}L_{1}}D)$ by a factor of $\sqrt{k/d}$, thereby decreasing sensitivity and reducing the required noise intensity. For each client’s sensitive local dataset $\zeta_{i}$, $\forall i\in[n]$, if we treat DP-FCRN as the algorithm $\mathcal{A}$ defined in Definition 2, the worst-case observation by the attacker $\mathcal{A}(\zeta_{i})=\{\mathcal{S}(y_{i,t})|0\leq t\leq T\}$. Theorem 1 states a sufficient condition for achieving $(\varepsilon,\delta)$-DP based on the reduced sensitivity resulting from sparsification.
@@
Suppose Assumption 1 holds, and the random-$k$ sparsifier with $k\leq d$ is used in Algorithm 1. Given $m$, $\tau$, $\varepsilon\in(0,1]$, and $\delta_{0}\in(0,1]$, if the noise variance
$\sigma^{2}\geq\frac{160\tau Tk\log(1.25/\delta_{0}){({\color[rgb]{0,0,1}L_{0}}+{\color[rgb]{0,0,1}L_{1}}D)}^{2}}{\varepsilon^{2}m^{2}d}$ (7)
and $T\geq\frac{\varepsilon^{2}}{4\tau}$, then DP-FCRN is $(\varepsilon,\delta)$-DP for $\zeta_{i}$, $\forall i\in[n]$, with some constant $\delta\in(0,1]$. Specifically, for any output set of DP-FCRN, $\mathcal{A}(\zeta_{i})$, we have
${\color[rgb]{0,0,1}\mathbb{P}\{\mathcal{A}(\zeta_{i})\in\mathcal{O}\}\leq e^{\varepsilon}\mathbb{P}\{\mathcal{A}(\zeta_{i}^{\prime})\in\mathcal{O}\}+\delta.}$ (8)
@@
The required noise intensity is proportional to the sparsification ratio, $k/d$. Therefore, to achieve the same level of DP, the required noise under our algorithm can be reduced by decreasing $k$. In other words, the fewer transmitted bits, the less noise required for $(\varepsilon,\delta)$-DP. The assumption that $\varepsilon\in(0,1]$ is motivated by the need to ensure the validity of theoretical results, such as composition theorems and privacy amplification, which often require $\varepsilon$ to be small. Additionally, small $\varepsilon$ aligns with the goal of providing strong privacy guarantees, making this range both theoretically and practically relevant for differential privacy research.
Remark 4.
Common compression methods include quantizers and sparsifiers. However, quantization can increase the sensitivity of gradient updates and disrupt the distribution of Gaussian perturbations, making both algorithm design and analysis more difficult. In contrast, sparsifiers simply set some elements to zero, reducing the sensitivity of the messages and making privacy amplification more tractable. Moreover, compared to the Top-$k$ sparsifier, the random-$k$ sparsifier introduces additional randomness, further enhancing the privacy guarantee. In the future, it will be interesting to study the potential privacy amplification under other compression schemes.
@@ 5 Convergence Analysis
This section presents the convergence analysis of Algorithm 1. In each step of the algorithm, a global cubic upper bound function $\phi:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ for $f(w)$ is constructed as
@@
Algorithm 2 uses the standard SGD to solve (5) and the local cubic upper bound $\phi_{i}(x;x_{t})$ is a strongly convex function. Since each sample $\zeta_{i}^{(j)}$ is chosen with equal probability, i.e., $\mathbb{P}(\zeta_{i}^{(j)})=1/m$, $\mathbb{E}[\hat{g}_{i,t}]=\sum_{j=1}^{m}\nabla l(x,\zeta_{i}^{(j)})\mathbb{P}(\zeta_{i}^{(j)})=\nabla f_{i}(x)$ and $\mathbb{E}[\hat{H}_{i,t}]=\sum_{j=1}^{m}\nabla^{2}l(x,\zeta_{i}^{(j)})\mathbb{P}(\zeta_{i}^{(j)})=\nabla^{2}f_{i}(x)$ are unbiased estimates. Therefore, we can obtain the suboptimality gap based on the typical SGD analysis.
@@
Suppose that Assumptions 21 hold. Given parameters $\varepsilon\in(0,1]$, $\delta_{0}\in(0,1]$, and $w\in\mathcal{X}$ the output of Algorithm 2, if we set the number of local iterations as
$\tau=\frac{(L_{0}+L_{1}D+MD^{2}/2)^{2}\varepsilon^{2}m^{2}}{kT\log(1/\delta_{0})({\color[rgb]{0,0,1}L_{0}}+{\color[rgb]{0,0,1}L_{1}}D)^{2}},$ (10)
@@
$=$ $O\left(\frac{k\log(1/\delta_{0}){({\color[rgb]{0,0,1}L_{0}}+{\color[rgb]{0,0,1}L_{1}}D)}^{2}T}{\varepsilon^{2}m^{2}\mu}\right).$
@@
Suppose that Assumptions 21 hold and the random-$k$ sparsifier with $k\leq d$ is used in Algorithm 1. Given parameters $m$ and $\varepsilon\in(0,1]$, $\delta_{0}\in(0,1]$, by setting the number of local iterations as (10), the step size as $\alpha>1$ and
$\alpha=O\left(\frac{k\log(1/\delta_{0}){({\color[rgb]{0,0,1}L_{0}}+{\color[rgb]{0,0,1}L_{1}}D)}^{2}T}{\varepsilon^{2}m^{2}\mu(L_{0}+L_{1}D+MD^{2}/2)D}\right),$
@@
$\leq$ $\tilde{O}\left(\frac{k\log(1/\delta_{0}){({\color[rgb]{0,0,1}L_{0}}+{\color[rgb]{0,0,1}L_{1}}D)}^{2}}{\varepsilon^{2}m^{2}\mu}\cdot\frac{\sqrt{L_{2}}(f(x_{0})-f(x^{*}))^{\frac{1}{4}}}{\mu^{\frac{3}{4}}}\right).$
@@
With the boundedness established in Assumption 2, existing DP algorithms for strongly convex functions achieve the best bound for optimization error, $O\left(\frac{d}{\varepsilon^{2}}\right)$ (Bassily et al. (2014); Kairouz et al. (2021)). This indicates that the error bound derived in Theorem 2 is optimal w.r.t. the privacy loss $\varepsilon$. Furthermore, our result $O\left(\frac{k}{\varepsilon^{2}}\right)$ reduces the error bound by a factor $k/d$, attributed to sparsification. This result underscores how efficient communication better balances the trade-off between privacy and utility. Unlike the recent algorithm in Chen et al. (2024), which assumes a trusted central server, we adopt a client-level differential privacy (DP) approach that offers stronger and more robust privacy protection. Furthermore, the error bound in Chen et al. (2024) increases when fewer coordinates are transmitted, implying that higher communication efficiency leads to worse convergence accuracy. In contrast, the error bound of our algorithm shows that more efficient communication reduces convergence error. Thus, in the context of federated second-order learning, we are the first to improve the trade-off between privacy and accuracy through efficient communication.
@@
While DP-FCRN does not explicitly include a switching step, the proof of Theorem 2 indicates that DP-FCRN operates in two distinct phases. Initially, when $x_{t}$ is distant from $x^{*}$, the convergence rate is $1/T^{4}$. Subsequently, as $x_{t}$ approaches $x^{*}$, the algorithm transitions to the second phase with a convergence rate of $\exp(\exp(-T))$. In summary, leveraging second-order techniques in our algorithm significantly improves the oracle complexity compared to first-order methods (Liu et al. (2024)).
Remark 7.
The privacy analysis is independent of convexity assumptions. While many non-convex algorithms focus on first-order stationary points, which may be poor local minima or saddle points, future work will explore convergence to second-order stationary points using cubic regularization. This can reduce the risk of saddle points and improve local minima. Additionally, time-varying step sizes will be essential for optimizing the achievable bounds.
@@ 6 Numerical Evaluation
In this section, we evaluate the effectiveness of DP-FCRN with different sparsification ratios and compare them to the first-order Fed-SGD with DP (Lowy & Razaviyayn (2023)).
@@ 6.1 Experimental Setup
We test our algorithm on the benchmark datasets epsilon (Sonnenburg et al. (2006)), which include 400,000 samples and 2,000 features for each sample. The data samples are evenly and randomly allocated among the $n=40$ clients. The clients cooperatively solve the following logistic regression problem:
@@
As DP parameters, we consider $\varepsilon\in\{0.4,0.6,0.8,1\}$ and $\delta_{0}=0.01$. The random noise is generated according to (7). We choose $\alpha=1$. As for Fed-SGD, we set the learning rate as one. Moreover, $L_{0}=0.1$, $L_{1}=1$, $M=1$, $D=0.1$, $\delta_{0}=0.01$, and we calculate the value of $\tau$ using (10). In iteration $t$, client $i$ processes one data point from $\zeta_{i}$ and the server updates $x_{t}$ accordingly. Upon finishing processing the entire dataset, one epoch is completed. We conduct the algorithm for four epochs and repeat each experiment five times. We show the mean curve along with the region representing one standard deviation. The convergence performance of the algorithm is evaluated by training suboptimality and testing accuracy over iterations. Training suboptimality is calculated by $f(x_{t})-f(x^{*})$, where $f(x^{*})$ is obtained using the LogisticSGD optimizer from scikit-learn (Pedregosa et al. (2011)). Testing accuracy is determined by applying the logistic function to the entire dataset. It is calculated as the percentage of correct predictions out of the total number of predictions.
@@ 6.2 Performance and Comparison with Fed-SGD
Moreover, employing a larger sparsification ratio $k/d$ in DP-FCRN results in worse training suboptimality, verifying Theorem 2. We find that keeping more coordinates in sparsification leads to more complete information transmission together with increased noise. The results shown in Fig. 1 indicate that, in certain settings, the benefit of noise reduction for convergence performance may outweigh the negative impacts arising from information completeness. On the other hand, there is no obvious difference in testing accuracy with different sparsification ratios, which indicates that the performance under the proposed DP-FCRN does not deteriorate much while reducing the communication burden.
@@ 7 Conclusion and Future Work
There are several promising directions for future research. Firstly, investigating methods to reduce the computational complexity of federated second-order learning approaches is valuable. Additionally, exploring communication-efficient and privacy-preserving variants of advanced federated second-order algorithms, such as GIANT and SHED, presents promising research directions.
@@ Appendix A Supporting Lemmas
We begin by introducing some properties of random-$k$ sparsification.
Lemma 3.
The following lemma provide some useful properties of $\phi_{i}(v;w)$ (Ganesh et al. (2024)).
@@
(Ganesh et al. (2024)) Let $b_{0}>0$ and define the sequence $a_{t+1}\leq b_{0}+\frac{1}{2}a_{t}^{\frac{3}{2}}$ where $a_{0}\leq\frac{16}{9}$. Then, after $T=\Theta(\log\log(\frac{1}{b}))$, we have $a_{T}=O(b_{0})$.
@@ Appendix B Proof of Theorem 1
(Privacy for Subsampling (Steinke (2022))) Suppose $\mathcal{G}$ is an $(\varepsilon,\delta)$-DP mechanism. Consider $\text{Sample}_{r_{1},r_{2}}:\mathcal{D}^{r_{1}}\to\mathcal{D}^{r_{2}}$ as the subsampling manipulation. Given a dataset belonging to $\mathcal{D}^{r_{1}}$ as an input, this subsampling manipulation selects a subset of $r_{2}\leq r_{1}$ elements from the input dataset uniformly at random. For the following mechanism
@@
(Composition of DP (Steinke (2022))) If each of $T$ randomized algorithms $\mathcal{A}_{1},\dots,\mathcal{A}_{T}$ is $(\varepsilon_{i},\delta_{i})$-DP with $\varepsilon_{i}\in(0,0.9]$ and $\delta_{i}\in(0,1]$, then $\mathcal{A}$ with $\mathcal{A}(\cdot)=(\mathcal{A}_{1}(\cdot),\dots,\mathcal{A}_{T}(\cdot))$ is $(\tilde{\varepsilon},\tilde{\delta})$-DP with
@@
$\varepsilon_{s}=\frac{2\sqrt{2k\log(1.25/\delta_{0})}{({\color[rgb]{0,0,1}L_{0}}+{\color[rgb]{0,0,1}L_{1}}D)}}{\sigma\sqrt{d}}$
@@
$\varepsilon_{s}^{\prime 2}\leq\frac{32k\log(1.25/\delta_{0})(L_{0}+L_{1}D)^{2}}{\sigma^{2}m^{2}d}\leq\frac{\varepsilon^{2}}{5\tau T}\leq 0.8.$
Therefore, we have $\varepsilon_{s}^{\prime}\leq 0.9$ and
$\sum_{s=1}^{\tau T}\varepsilon_{s}^{\prime 2}\leq\frac{1}{5}\sum_{s=1}^{\tau T}\frac{\varepsilon^{2}}{\tau T}\leq 1$ (14)
@@
$\tilde{\varepsilon}=\sqrt{\sum_{s=1}^{\tau T}2\varepsilon_{s}^{\prime 2}\log\left(e+\frac{\sqrt{\sum_{s=1}^{\tau T}\varepsilon_{s}^{\prime 2}}}{\tilde{\delta}}\right)}+\sum_{s=1}^{\tau T}\varepsilon_{s}^{\prime 2}$
@@
$\tilde{\varepsilon}\leq$ $\sqrt{\sum_{s=1}^{\tau T}2\varepsilon_{s}^{\prime 2}\log\left(e+\frac{\sqrt{\sum_{s=1}^{\tau T}\varepsilon_{s}^{\prime 2}}}{\tilde{\delta}}\right)}+\frac{1}{5}\varepsilon^{2}$
$\leq$ $\sqrt{3\sum_{s=1}^{\tau T}\varepsilon_{s}^{\prime 2}}+\frac{1}{5}\varepsilon$